OpenStudy (anonymous):

with slope -4 that passes through 0,-5

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-(-5)=(-4)(x-0)\] \[y+5=-4x\] \[y=-4x-5\]